OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1

OpenStudy (anonymous):

@cram

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

I have ten minutes left on this test, and I have three questions to ask on here :)

OpenStudy (texaschic101):

you have to first try to get a variable by itself. 2x + y = 1 (subtract 2x from both sides) y = -2x + 1 now sub -2x + 1 in for y in the other equation.. 4x + 2y = -1 4x + 2(-2x + 1) = -1 4x - 4x + 2 = -1 2 = -1 (incorrect) this problem has NO SOLUTION

OpenStudy (anonymous):

Can you help me with a couple more?

OpenStudy (texaschic101):

ask away...you don't have much time

OpenStudy (anonymous):

Determine which system below will produce infinitely many solutions. 2x + 5y = 24 2x + 5y = 42

OpenStudy (anonymous):

2x + 5y = 24 2x + 5y = 42 3x - 2y = 15 6x + 5y = 11 4x - 3y = 9 -8x + 6y = -18 5x - 3y = 16 -2x + 3y = -7

OpenStudy (texaschic101):

-8x + 6y = -18 ---- reduce by dividing by 2 -4x + 3y = -9 -- multiply by -1 to make x positive 4x - 3y = 9 same as other equation...so infinite solutions answer is 3rd one

OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18 12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = -15 3y = 9

OpenStudy (texaschic101):

crap...I read this wrong...do you mean the same solution as the first one ?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Its the same kind of problem you did before

OpenStudy (anonymous):

I had to submit it

OpenStudy (texaschic101):

no its not....produce the same solution as what ? Oh...I am so sorry.

OpenStudy (anonymous):

Wait its letting me do a question? lol.. Whatever its this one Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18 12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = -15 3y = 9

OpenStudy (anonymous):

Oh its the same quetion.. Well its letting me do it again, just this question to improve my score. so lets go for it!

OpenStudy (texaschic101):

6x + 2y = -6 -- reduce 3x + y = - 3 y = -3x - 3 3x - 4y = -18 3x - 4(-3x - 3) = -18 3x + 12x + 12 = -18 15x + 12 = -18 15x = -18 - 12 15x = -30 x = -2 y = -3x - 3 y = -3(-2) - 3 y = 6 - 3 y = 3 solution is (-2,3) ----------------------- 12x + 4y = -12 -- reduce 3x + y = -3 y = -3x - 3 15x = -30 x = -2 y = -3x - 3 y = -3(-2) - 3 y = 6 - 3 y = 3 solution is (-2,3) --------------------- your first two sets of equations have the same solution

OpenStudy (anonymous):

so what answer choice?

OpenStudy (texaschic101):

I think the question was asking what has the same solution as the first set of equations....if that is the case, then the answer is : 12x + 4y = -12....15x = -30, because the first one is what you are comparing the answer to

OpenStudy (anonymous):

Ah ok :)

OpenStudy (texaschic101):

sorry it took so long

OpenStudy (anonymous):

I got a 90! :)

OpenStudy (texaschic101):

WHOOP - WHOOP...doing the happy dance now. Way to go :)

OpenStudy (anonymous):

haha I am going to be doing a ton of assignments because I thought my class ended 9/8, not 8/8 and I have 20 assignments left. can you help me? I am doing 10 today. now 9. Can you help with a few atleast?

OpenStudy (texaschic101):

are they math ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

close this post and ask your questions

OpenStudy (texaschic101):

we can still keep typing even though this question is closed

OpenStudy (anonymous):

Alright I closed it

OpenStudy (anonymous):

Which of the following inequalities matches the graph? graph of an inequality with a solid horizontal line through the origin and shading below the line x less than or greater to 0 x greater than or equal to 0 y less than or greater to 0 y greater than or equal to 0

OpenStudy (texaschic101):

solid line means it has an equal sign.....horizontal line through the origin means x will always be 0. Shading below the line means less then. answer : y < = 0 ....its gonna be answer 3.....it should read less then or equal, not greater.

OpenStudy (anonymous):

Which correctly describes how the graph of the inequality -4y - x ≥ 7 is shaded? Above the solid line Below the solid line Above the dashed line Below the dashed line

OpenStudy (texaschic101):

-4y - x >= 7 -4y >= x + 7 -- when dividing by negative, the sign changes y <= -1/4x - 7/4 if there is an equal sign, it will be a solid line < less then means shading below the line so answer is : Below solid line

OpenStudy (anonymous):

Which of the following inequalities matches the graph? graph of an inequality with a solid line through the points (0, -2) and (2, 1) with shading above the line 3x - 2y greater than or equal to 4 3x - 4y less than or greater to 2 3x - 2y less than or greater to 4 The correct inequality is not listed.

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) slope = (1 - (-2) / (2 - 0) slope = 3/2 y = mx + b (2,1)...x = 2 and y = 1 slope = 3/2 1 = 3/2(2) + b 1 = 3 + b 1 - 3 = b -2 = b y = 3/2x - 2 --- multiply by 2 2y = 3x - 2 -- subtract 3x from both sides -3x + 2y = -2 -- multiply by -1 to make x positive 3x - 2y = 2 solid line means there is an equal sign. Shading above the line means greater then. answer is : Correct one is not listed. It should be 3x - 2y >= 2

OpenStudy (anonymous):

Which inequality matches the graph? X, Y graph. X range is negative 10 to 10, and Y range is negative 10 to 10. Dashed line on graph has positive slope and runs through negative 10, negative 9 and negative 1, negative 3 and 8, 3. Above line is shaded. -2x + 3y > 7 2x - 3y < 7 -3x + 2y ≥ 7 3x - 2y ≤ 7

OpenStudy (texaschic101):

Dashed line does not contain an equal sign...so last 2 answers are eliminated. -2x + 3y > 7 3y > 2x + 7 y > 2/3x + 7/3 positive slope -2x + 3y < 7 3y < 2x + 7 y < 2/3x + 7/3 positive slope above line shading means greater.....so answer is : first one

OpenStudy (anonymous):

Which of the following inequalities matches the graph? graph of an inequality with a solid vertical line through the point (5, 0) and shading to the left of the line x less than or greater to 5 x greater than or equal to 5 y less than or greater to 5 y greater than or equal to 5

OpenStudy (texaschic101):

solid vertical line passing through (5,0)....this means x willl always be 5. I am having trouble with this one....I am sorry

OpenStudy (anonymous):

kk next one

OpenStudy (texaschic101):

I know its gonna be less then

OpenStudy (anonymous):

Lincoln Middle School has a total of $200 to spend on notebooks and tablets for the school board meeting. Notebooks cost $7.00 each, and tablets cost $5.00 each. If x represents the number of notebooks and y represents the number of tablets, which graph correctly shows the solution to the problem? X, Y graph. X range is negative 60 to 60, and Y range is negative 60 to 60. Solid line on graph has negative slope and runs through 0, 40 and 50, negative 30. Below line is shaded. X, Y graph. X range is negative 60 to 44, and Y range is negative 60 to 60. Solid line on graph has negative slope and runs through 0, 40. Above line is shaded. X, Y graph. X range is negative 60 to 44, and Y range is negative 60 to 60. Dashed line on graph has negative slope and runs through 0, 40. Above line is shaded. X, Y graph. X range is negative 60 to 44, and Y range is negative 60 to 60. Dashed line on graph has negative slope and runs through 0, 40. Below line is shaded.

OpenStudy (texaschic101):

I think it will be x <= 5

OpenStudy (texaschic101):

thats for the last question

OpenStudy (texaschic101):

7x + 5y < = 200 5y <= -7x + 200 y <= -7/5x + 40 It contains a solid line because of the equal sign It has a negative slope y intercept is (0,40) answer is : I think it is the first one

OpenStudy (texaschic101):

I am so sorry but I have to go now. I hope I was able to help some. Good-bye for now.

OpenStudy (anonymous):

Wait! Theres on emore problem!

OpenStudy (texaschic101):

lets have it

OpenStudy (anonymous):

The networking organization you joined is throwing a party. You are in charge of buying the chips, which cost $2.50 per bag and salsa, which costs $4.00 per jar. The chips & salsa budget you are given totals $60. The inequality 2.5x + 4y 60 represents the possible combinations of chips (x) and salsa (y) you can buy. graph of the inequality 2.5 x plus 4 y is less than or equal to 60 Which of the following does not represent a solution to the inequality? 10 bags of chips and 2 jars of salsa 20 bags of chips and 2 jars of salsa 14 bags of chips and 5 jars of salsa 18 bags of chips and 6 jars of salsa

OpenStudy (anonymous):

When will you be back? lol you are like my hero.

OpenStudy (texaschic101):

2.5x + 4y <= 60 x = chips and salsa = y now we sub in our answer choices and see which ones are correct. 2.5(10) + 4(2) <= 60 25 + 8 <= 60 33 <= 60 (correct) 2.5(20) + 4(2) <= 60 50 + 8 <= 60 58 <= 60 (correct) 2.5(14) + 4(5) <= 60 35 + 20 <= 60 55 <= 60 (correct) 2.5(18) + 4(6) <= 60 45 + 24 <= 60 69 <= 60 (incorrect) answer is the last one. That figures....I had to do all the others to find this out...lol I will probably be back on tonight sometime...sorry that I can't give you a definite time.

OpenStudy (texaschic101):

I have to go now, but I will hopefully catch you later

OpenStudy (anonymous):

I got a 70 By the way. thatll have to do! Ill do the three assignments on my own and hopefully you can make a dent with me in the night assignments!